# Linear Preference Optimization: Decoupled Gradient Control via Absolute Regularization

### v1

@@ Abstract
DPO (Direct Preference Optimization) has become a widely used offline preference optimization algorithm due to its simplicity and training stability. However, DPO is prone to overfitting and collapse. To address these challenges, we propose Linear Preference Optimization (LPO), a novel alignment framework featuring three key innovations. First, we introduce gradient decoupling by replacing the log-sigmoid function with an absolute difference loss, thereby isolating the optimization dynamics. Second, we improve stability through an offset constraint combined with a positive regularization term to preserve the chosen response quality. Third, we implement controllable rejection suppression using gradient separation with straightforward estimation and a tunable coefficient that linearly regulates the descent of the rejection probability. Through extensive experiments, we demonstrate that LPO consistently improves performance on various tasks, including general text tasks, math tasks, and text-to-speech (TTS) tasks. These results establish LPO as a robust and tunable paradigm for preference alignment, and we release the source code, models, and training data publicly.
@@ 1 Introduction
Despite its elegance and widespread adoption, DPO exhibits several critical shortcomings. First, the inherent coupling within the logsigmoid function forces the optimization of the chosen response’s log-probability $\log\pi_{\theta}(y_{w}|x)$ and the rejected response’s log-probability $\log\pi_{\theta}(y_{l}|x)$ to be interdependent. This often manifests itself as an undesirable and significant decrease in the logarithmic probability of the responses chosen during training, which can degrade their inherent quality even as the preference objective improves. Second, DPO is highly sensitive to the quality and noise level within preference datasets. Suboptimal or ambiguous preference pairs can lead to overfitting and subpar performance. Third, DPO lacks explicit mechanisms to control the magnitude of the gap between the logarithmic probabilities of the chosen and rejected responses $\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$, which can lead to overoptimization and reduced generalization.
@@
  • Gradient Decoupling via Absolute Regulation: We replace the ’logsigmoid’ function with the absolute difference function. This crucial modification decouples the gradients flowing back to the chosen and rejected log-probabilities, enabling more independent and targeted optimization of each term.
  • Stability Enhancement via Offset and Positive Constraint: Inspired by Offset-DPO Amini et al. (2024) and Identity Preference Optimization (IPO) Azar et al. (2023), we introduce an offset $\mu$ to explicitly constrain the gap $\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$, preventing it from growing excessively large and improving generalization. Simultaneously, drawing motivation from DPOP Pal et al. (2024a), we incorporate an explicit positive constraint term $\log\pi_{\theta}(y_{w}|x)$ to counteract the problematic decrease in the log-probability of chosen responses observed in standard DPO.
  • Controlled Rejection Suppression via Gradient Separation: Leveraging the Straight-Through Estimator (STE) technique Esser et al. (2021), we strategically detach the computational graph (using ‘tensor.detach()‘) to isolate the gradients of the chosen and rejected log-probabilities. This allows us to introduce a control coefficient $r_{2}$ specifically on the gradient path influencing the rejected response’s log-probability. By modulating $r_{2}$, we gain fine-grained control over the rate at which the log-probability of rejected responses is suppressed during optimization.
We conduct extensive experiments to validate the effectiveness of LPO. Our results demonstrate:
@@
  • Specialized Task Superiority: LPO excels in specialized domains. Notably, models fine-tuned with LPO significantly outperform strong baselines like Qwen2.5-Instruct on mathematical reasoning (GSM8K Cobbe et al. (2021)). Furthermore, LPO yields substantial performance gains in specialized speech processing tasks, including Automatic Speech Recognition (ASR) and Text-to-Speech (TTS) synthesis within a native speech multimodality framework.
@@
  • We demonstrate and validate the practical utility of the $r_{2}$ coefficient for controlling rejection suppression rates.
@@ 2 Related Works
Current large language models (LLMs) demonstrate strong capability in following human instructionsYang et al. (2025); Liu et al. (2024). Reinforcement Learning from Human Feedback (RLHF)Schulman et al. (2017b) is a crucial technique for aligning these models with human values. However, RLHF training requires separately training a reward modelChristiano et al. (2017); Ouyang et al. (2022b), necessitating the simultaneous loading of four distinct models during the process. This approach consumes significant computational resources, and the training process can be unstableRafailov et al. (2023). To address these limitations, Direct Preference Optimization (DPO)Rafailov et al. (2023) proposes a novel parameterization method for the reward model. This solution enables derivation of the optimal policy via a closed-form solution, thereby allowing conventional RLHF problems to be solved using a simple classification loss function.
The DPO training process is prone to overfitting, with both positive and negative sample probabilities tending to decreaseFeng et al. (2024b). To address these issues, several improvements have been proposed, such as DPOPPal et al. (2024b) and IPOAzar et al. (2024). IPO analyzes the theoretical foundations of RLHF (Reinforcement Learning from Human Feedback) and DPO (Direct Preference Optimization). To counter DPO’s susceptibility to overfitting, it introduces a pairwise preference loss function termed ”Identity Preference Optimization (IPO)”. This method prevents overfitting on preference data by penalizing preference margins exceeding a regularization threshold. DPOPPal et al. (2024b) addresses the issue of declining positive sample probability by incorporating a penalty term for positive samples into its objective function. SimPOMeng et al. (2024) employs the average log probability of a sequence as its implicit reward function. This reward design enables more precise alignment of model generation behavior. Furthermore, by eliminating the dependency on a reference model, it significantly improves computational efficiency and reduces memory consumption. Our approach replaces DPO’s log-sigmoid function with an absolute function, incorporates SimPO’s length normalization, and decouples the gradient computations for positive and negative samples. This decoupling allows for explicit control over the magnitude of negative sample gradient descent.
@@ 3.1 Limitations of DPO
Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022a) has been demonstrated to significantly enhance the robustness of model responses, safety, and reasoning capabilities of large language models in previous training paradigms. However, the practical application of RLHF presents several challenges. These include the difficulty in training a sufficiently robust reward model to prevent reward hacking, while certain RLHF algorithms, such as Proximal Policy Optimization (PPO) Schulman et al. (2017a), suffer from difficult parameter tuning. Consequently, to streamline the alignment training process for large models, Direct Preference Optimization (DPO) Rafailov et al. (2024) has been proposed. DPO demonstrates that the RLHF training procedure can be simplified into a maximum likelihood optimization problem without the need to explicitly train a separate reward model:
$\begin{split}L_{\text{DPO}}(\pi_{\theta},\pi_{\text{ref}})=&-E_{(x,y_{w},y_{l})\sim D}\bigg{[}\log\sigma\bigg{(}\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\bigg{)}\bigg{]}\end{split}$ (1)
Among Eq.1, $\sigma$ denotes the logistic function; $D=\{(x^{i},y_{w}^{i},y_{l}^{i})\}_{i=1}^{N}$, $x^{i}$ represents the prompt, and $y_{w}^{i}$ and $y_{l}^{i}$ represent the chosen response and rejected response under the input prompt $x$ respectively; $\pi_{\theta}$ denotes the policy model to be optimized, initialized from the SFT model, and $\pi_{\text{ref}}$ denotes the Reference model, which is the SFT model.
@@
Among Eq.2, $x_{1}$ represents $\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}$, and $x_{2}$ represents $\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$. Then, we take the partial derivatives with respect to $x_{1}$ and $x_{2}$ respectively, and we can obtain:
@@
Among Eq.3, we divide $\frac{\partial L_{DPO}(x_{1},x_{2})}{\partial x_{1}}$ by $\frac{\partial L_{DPO}(x_{1},x_{2})}{\partial x_{2}}$, and we can obtain the following:
@@
Based on the characteristics of the BT model and the DPO training objectives, we can conclude that if we want to maximize the likelihood function of the DPO, then $x_{1}-x_{2}$ must be positive, which means $x_{1}>x_{2}$. Therefore, the gradient generated by $x_{1}$ (corresponding to chosen) will be less than that generated by $x_{2}$ (corresponding to reject). Also, due to the marginal effect of the logistic function, in the later stages of training, $x_{2}$ will be much smaller than $x_{1}$, so the gradient generated by $x_{2}$ will dominate in the later stages of training. This means that the logprobs of rejection will be reduced to an extremely low value. However, in the actual training of DPO, there is no need to reduce the logprobs of reject so much.
@@
Among the three cases, Case 1 represents the ideal optimization target for DPO, where $x_{1}$ slightly increases while $x_{2}$ decreases at a reasonable rate. However, in practical DPO training, there is often a problem in which $x_{1}$ and $x_{2}$ decrease simultaneously, leading to a decrease in model performance.
Summarizing the limitations of DPO in two aspects:
(i) The contribution of the chosen to the gradient is always less than that of the rejection, causing the optimization target to focus on reducing the logprobs of the rejection. However, due to the nature of the sigmoid function, the reduction in reject logprobs becomes excessively large.
(ii) Since the objective of DPO training is essentially to increase the difference between $\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}$ and $\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$, it often results in both decreasing simultaneously, which causes the performance of the model to decline rather than improve.
@@ 3.2 Linear Preference Optimization: Decoupling the grad between chosen and reject
Eq.1 shows that the DPO target function can be abstracted as $L_{DPO}(x_{1},x_{2})=f(x_{1},x_{2})$. From Eq.2, the gradients $\frac{\partial L_{DPO}(x_{1},x_{2})}{\partial x_{1}}$ and $\frac{\partial L_{DPO}(x_{1},x_{2})}{\partial x_{2}}$ contain nonlinear terms of $x_{1}$ and $x_{2}$. Hence, we linearize the mathematical expression of DPO.
Replace DPO’s logsigmoid function with the abs function. Add an Offset inspired by IPO Azar et al. (2023) and ODPO Amini et al. (2024), include a Positive term inspired by DPOP Pal et al. (2024a), and apply length normalization to chosen and reject logprobs as in SimPO Meng et al. (2024). The Linear Preference Optimization function becomes:
$\begin{split}\mathcal{L}_{LPO}=2\beta\left|x_{1}-x_{2}-\frac{1}{2\beta}\right|+\lambda\cdot\max(0,-x_{1})\end{split}$ (6)
Among Eq.6, hyperparameter $\beta$ controls the offset, $\lambda$ controls the magnitude of the positive term, and $x_{1}$, $x_{2}$ are the length-normalized logprobs:
@@
Among Eq.7, $len_{w}$ denotes the length of the chosen response, and $len_{l}$ denotes the length of the reject response.
Using the gradient analysis method Feng et al. (2024a), compute the partial derivatives with respect to $x_{1}$ and $x_{2}$:
$\begin{split}\begin{cases}\frac{\partial L_{LPO}(x_{1},x_{2})}{\partial x_{1}}=-2\beta\cdot sgn(x_{1}-x_{2}-\frac{1}{2\beta})+C\\[12.91663pt] \frac{\partial L_{LPO}(x_{1},x_{2})}{\partial x_{2}}=-2\beta\cdot sgn(x_{1}-x_{2}-\frac{1}{2\beta})\end{cases}\end{split}$ (8)
Where $sgn(u)$ is the sign function, which is 1 if $u>0$, -1 if $u<0$. the expression for C is also a constant, as shown in Eq9
@@
We divide $\frac{\partial L_{LPO}\left(x_{1},x_{2}\right)}{\partial x_{1}}$ by $\frac{\partial L_{LPO}\left(x_{1},x_{2}\right)}{\partial x_{2}}$ and obtain the following:
$\begin{split}\frac{\partial L_{LPO}\left(x_{1},x_{2}\right)}{\partial x_{1}}\bigg{/}\frac{\partial L_{LPO}\left(x_{1},x_{2}\right)}{\partial x_{2}}=\frac{-2\beta\cdot sgn(x_{1}-x_{2}-\frac{1}{2\beta})+C}{2\beta\cdot sgn(x_{1}-x_{2}-\frac{1}{2\beta})}\end{split}$ (10)
From Eq.10, it can be seen that the gradient ratio of $x_{1}$ and $x_{2}$ becomes a constant, and the relative magnitude of their gradients can be controlled through $\beta$ and $\lambda$. Meanwhile, in order to better control the descent rate of $x_{2}$, we use the straight-through estimator (STE) Esser et al. (2021) to improve Eq.1, resulting in LPO-ste:
$\begin{split}\begin{cases}L_{LPO-ste}^{x_{1}}=r_{1}\cdot 2\beta\left|x_{1}-x_{2}.detach()-\frac{1}{2\beta}\right|+\lambda\cdot\max(0,-x_{1})\\[8.61108pt] L_{LPO-ste}^{x_{2}}=r_{2}\cdot 2\beta\left|x_{1}.detach()-x_{2}-\frac{1}{2\beta}\right|+\lambda\cdot\max(0,-x_{1}.detach())\end{cases}\end{split}$ (11)
By using STE, the gradients of $L_{LPO-ste}^{x_{1}}$ and $L_{LPO-ste}^{x_{2}}$ in Eq.11 are fully separated. Thus, the descent rates of $x_{1}$ and $x_{2}$ can be controlled by $r_{1}$ and $r_{2}$, respectively. This achieves both gradient separation and descent rate control.
Ultimately, the expression for LPO-ste is:
@@
In the practical application of LPOsteLPO-ste, $r_{1}$ is generally fixed at 1.0, while $r_{2}$ is adjusted within the range $[0.05,3.0]$. Fig 1 shows the descent rate of reject and the ascent rate of chosen under different $r_{2}$ values. From Fig 1, we can observe that adjusting the size of $r_{2}$ allows for control over the ascent rate of chosen and the descent rate of reject, thereby achieving the goal of adjusting the model’s performance.

### v2

@@ Abstract
DPO (Direct Preference Optimization) has become a widely used offline preference optimization algorithm due to its simplicity and training stability. However, DPO is prone to overfitting and collapse. To address these challenges, we propose Linear Preference Optimization (LPO), a novel alignment framework featuring three key innovations. First, we introduce gradient decoupling by replacing the log-sigmoid function with an absolute difference loss, thereby isolating the optimization dynamics. Second, we improve stability through an offset constraint combined with a positive regularization term to preserve the chosen response quality. Third, we implement controllable rejection suppression using gradient separation with straightforward estimation and a tunable coefficient that linearly regulates the descent of the rejection probability. Through extensive experiments, we demonstrate that LPO consistently improves performance on various tasks, including general text, math, text-to-speech (TTS), and automatic speech recognition (ASR) tasks. These results establish LPO as a robust and tunable paradigm for preference alignment. Source code and model checkpoints are publicly available at https://github.com/IDEA-Emdoor-Lab/LPO.
@@ 1 Introduction
Despite its elegance and widespread adoption, DPO exhibits several critical shortcomings. First, the inherent coupling within the log-sigmoid function forces the optimization of log-probability of the chosen $\log\pi_{\theta}(y_{w}|x)$ and the rejected response $\log\pi_{\theta}(y_{l}|x)$ to be interdependent. This often manifests itself as an undesirable and significant decrease in the logarithmic probability of the responses chosen during training, which can degrade their inherent quality even as the preference objective improves. Second, DPO is highly sensitive to the quality and noise level within preference datasets. Suboptimal or ambiguous preference pairs can lead to overfitting and subpar performance. Third, DPO lacks explicit mechanisms to control the magnitude of the gap between the logarithmic probabilities of the chosen and rejected responses $\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$, which can lead to overoptimization and reduced generalization.
@@
  • Gradient Decoupling via Absolute Regulation: We replace the ’log-sigmoid’ function with the absolute difference function. This crucial modification decouples the gradients flowing back to the chosen and rejected log-probabilities, enabling more independent and targeted optimization of each term.
  • Stability Enhancement via Offset and Positive Constraint: Inspired by Offset-DPO (ODPO) Amini et al. (2024) and Identity Preference Optimization (IPO) Azar et al. (2023), we introduce an offset $\mu$ to explicitly constrain the gap $\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$, preventing it from growing excessively large and improving generalization. Simultaneously, inspired by DPOP Pal et al. (2024a), we incorporate an explicit positive constraint term $\log\pi_{\theta}(y_{w}|x)$ to counteract the problematic decrease in log-probability of chosen responses observed in standard DPO.
  • Controlled Rejection Suppression via Gradient Separation: Leveraging the Straight-Through Estimator (STE) technique Esser et al. (2021), we strategically detach the computational graph (using ‘tensor.detach()‘) to isolate the gradients of the chosen and rejected log-probabilities. This allows us to introduce a control coefficient $r_{2}$ specifically on the gradient path influencing the log-probability of the rejected response. By modulating $r_{2}$, we gain fine-grained control over the rate at which the log-probability of rejected responses is suppressed during optimization.
We conducted extensive experiments to validate the effectiveness of LPO. Our results demonstrate:
@@
  • Specialized Task Superiority: LPO excels in specialized domains. In particular, models fine-tuned with LPO significantly outperform strong baselines such as Qwen2.5-Instruct on mathematical reasoning (GSM8K Cobbe et al. (2021)). Furthermore, LPO yields substantial performance gains in specialized speech processing tasks, including Automatic Speech Recognition (ASR) and Text-to-Speech (TTS) synthesis within a native speech multimodality framework.
@@
  • We demonstrate and validate the practical utility of the coefficient $r_{2}$ for controlling rejection suppression rates.
@@ 2 Related Works
Current large language models (LLMs) demonstrate a strong capability in following human instructions Yang et al. (2025); Liu et al. (2024), showcasing their effectiveness in various applications such as text generation, question answering, and conversational agents. These models leverage extensive training on diverse datasets, allowing them to understand and respond accurately to user inputs. Their performance is further enhanced through techniques like Reinforcement Learning from Human Feedback (RLHF), which aligns model outputs with human preferences Schulman et al. (2017b). This continuous refinement of LLMs not only improves their responsiveness but also ensures they adhere to ethical guidelines and user expectations. However, RLHF training necessitates the separate training of a reward model Christiano et al. (2017); Ouyang et al. (2022b), which in turn requires the simultaneous loading of four distinct models. This method can be resource-intensive and may lead to instability during the training process Rafailov et al. (2023). To mitigate these issues, Direct Preference Optimization (DPO, Rafailov et al. (2023)) introduces a novel parameterization method for the reward model. This innovative approach facilitates the derivation of the optimal policy through a closed-form solution, enabling traditional RLHF challenges to be addressed using a straightforward classification loss function.
The DPO training process is prone to overfitting, with both positive and negative sample probabilities tending to decrease Feng et al. (2024b). To address these issues, several improvements have been proposed, such as DPOP Pal et al. (2024b) and IPO Azar et al. (2024). IPO examines the theoretical underpinnings of RLHF and DPO. To address DPO’s tendency to overfit, IPO introduces a pairwise preference loss function known as “Identity Preference Optimization (IPO).” This function is designed to mitigate overfitting on preference data by penalizing preference margins that exceed a specified regularization threshold, thereby enhancing the model’s generalization capabilities. DPOP Pal et al. (2024b) addresses the issue of declining positive sample probability by incorporating a penalty term for positive samples into its objective function. SimPO Meng et al. (2024) utilizes the average log-probability of a sequence as its implicit reward function, which allows for more accurate alignment of the model’s generation behavior. This reward structure not only enhances the effectiveness of the model but also eliminates the need for a reference model. As a result, SimPO significantly boosts computational efficiency and minimizes memory usage, making it a more resource-friendly approach in preference optimization. Our approach modifies DPO by replacing the log-sigmoid function with an absolute function. We also integrate SimPO’s length normalization and decouple the gradient computations for positive and negative samples. This decoupling provides explicit control over the magnitude of gradient descent for negative samples, enabling more targeted optimization and improving the overall performance of the model.
@@ 3.1 Limitations of DPO
RLHF Ouyang et al. (2022a) has been demonstrated to significantly enhance the robustness of model responses, safety, and reasoning capabilities of large language models in previous training paradigms. However, the practical application of RLHF presents several challenges. These include the difficulty in training a sufficiently robust reward model to prevent reward hacking, while certain RLHF algorithms, such as Proximal Policy Optimization (PPO) Schulman et al. (2017a), suffer from difficult parameter tuning. Consequently, to streamline the alignment training process for large models, Direct Preference Optimization (DPO) Rafailov et al. (2024) has been proposed. DPO demonstrates that the RLHF training procedure can be simplified into a maximum likelihood optimization problem without the need to explicitly train a separate reward model:
$\begin{split}L_{\text{DPO}}(\pi_{\theta},\pi_{\text{ref}})=&-E_{(x,y_{w},y_{l})\sim D}\bigg{[}\log\sigma\bigg{(}\beta\cdot\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\cdot\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\bigg{)}\bigg{]}\end{split}$ (1)
In Eq.1, $\sigma$ signifies the logistic function; $D=\{(x^{i},y_{w}^{i},y_{l}^{i})\}_{i=1}^{N}$ represents the dataset, where $x^{i}$ represents the prompt, and $y_{w}^{i}$ and $y_{l}^{i}$ denote the chosen response and rejected response for the input prompt $x$ respectively; The term $\pi_{\theta}$ refers to the policy model to be optimized, which is initialized from the Supervised Fine-Tuning (SFT) model, while $\pi_{\text{ref}}$ denotes the Reference model, also derived from the SFT model.
@@
In Eq.2, $x_{1}$ represents $\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}$, and $x_{2}$ represents $\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$. Then, we take the partial derivatives with respect to $x_{1}$ and $x_{2}$ respectively, and we can obtain:
@@
In Eq.3, $\frac{\partial L_{DPO}(x_{1},x_{2})}{\partial x_{1}}$ is divided by $\frac{\partial L_{DPO}(x_{1},x_{2})}{\partial x_{2}}$, and then we can obtain the following:
@@
Based on the characteristics of the BT model and the DPO training objectives, we can conclude that to maximize the likelihood function of the DPO, the condition $x_{1}-x_{2}$ must be positive, implying that $x_{1}>x_{2}$. Consequently, the gradient produced by $x_{1}$ (corresponding to the chosen response) will be less than that generated by $x_{2}$ (corresponding to the rejected response). Additionally, due to the marginal effect of the logistic function, in the later stages of training, $x_{2}$ will become significantly smaller than $x_{1}$, leading to the gradient from $x_{2}$ dominating the training process. This results in the log-probabilities of the rejected responses being reduced to extremely low values. However, in practice, there is no necessity to decrease the log-probabilities of the rejected responses to such an extent.
@@
Among the three cases, Case 1 represents the ideal optimization target for DPO, where $x_{1}$ slightly increases while $x_{2}$ decreases at a reasonable rate. However, in practical DPO training, a common issue arises where both $x_{1}$ and $x_{2}$ decrease simultaneously. This simultaneous decline can lead to a reduction in model performance, undermining the effectiveness of the training process.
The limitations of DPO can be summarized in two key aspects:
(i) The contribution of the chosen responses to the gradient is consistently less than that of the rejected responses. This imbalance results in the optimization process predominantly focusing on reducing the log-probabilities of the rejected samples. Consequently, the nature of the sigmoid function exacerbates this issue, leading to an excessively large reduction in the log-probabilities of the rejected responses.
(ii) Since the objective of DPO training is fundamentally to increase the difference between $\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}$ and $\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$, it frequently leads to both values decreasing simultaneously. This simultaneous decline results in a reduction in the model’s performance rather than an improvement, undermining the effectiveness of the training process.
@@ 3.2 Linear Preference Optimization: Decoupling the grad between chosen and reject
Eq.1 shows that the DPO target function can be represented as $L_{DPO}(x_{1},x_{2})=f(x_{1},x_{2})$. From Eq.2, the gradients $\frac{\partial L_{DPO}(x_{1},x_{2})}{\partial x_{1}}$ and $\frac{\partial L_{DPO}(x_{1},x_{2})}{\partial x_{2}}$ incorporate nonlinear terms involving both $x_{1}$ and $x_{2}$. Therefore, we proceed to linearize the mathematical expression of DPO to facilitate further analysis and optimization.
To enhance LPO function, we replace DPO’s log-sigmoid function with the absolute function. We also introduce an offset inspired by IPO Azar et al. (2023) and ODPO Amini et al. (2024), incorporate a positive term motivated by DPOP Pal et al. (2024a), and apply length normalization to both chosen and rejected log-probabilities similar to SimPO Meng et al. (2024). The resulting LPO function can be expressed as follows:
$\begin{split}\mathcal{L}_{LPO}=2\beta\cdot\left|x_{1}-x_{2}-\frac{1}{2\beta}\right|+\lambda\cdot\max(0,-x_{1})\end{split}$ (6)
Here, $\beta$ and $\lambda$ are hyperparameters controlling the offset and the magnitude of the positive term, respectively, while $x_{1}$ and $x_{2}$ represent the length-normalized log-probabilities of the chosen and rejected responses:
@@
Where $len_{w}$ denotes the length of the chosen response, and $len_{l}$ denotes the length of the rejected response.
Using the gradient analysis method Feng et al. (2024a), we can compute the partial derivatives of LPO function with respect to the variables $x_{1}$ and $x_{2}$. The derivatives can be expressed as follows:
$\begin{split}\begin{cases}\frac{\partial L_{LPO}(x_{1},x_{2})}{\partial x_{1}}=-2\beta\cdot\text{sgn}(x_{1}-x_{2}-\frac{1}{2\beta})+C\\[12.91663pt] \frac{\partial L_{LPO}(x_{1},x_{2})}{\partial x_{2}}=-2\beta\cdot\text{sgn}(x_{1}-x_{2}-\frac{1}{2\beta})\end{cases}\end{split}$ (8)
Where sgn($u$) is the sign function, which is defined as 1 if $u>0$, -1 if $u<0$, and 0 if $u=0$. The expression for $C$ is also a constant, as shown in Eq. 9. This constant plays a crucial role in adjusting the gradient based on the value of $x_{1}$, ensuring that the optimization process is influenced appropriately depending on whether the chosen response log-probability is negative or not. These partial derivatives allow us to analyze how changes in the log-probabilities of the chosen and rejected responses affect the overall optimization objective, providing insights into the dynamics of the optimization process.
@@
We divide $\frac{\partial L_{LPO}\left(x_{1},x_{2}\right)}{\partial x_{1}}$ by $\frac{\partial L_{LPO}\left(x_{1},x_{2}\right)}{\partial x_{2}}$ to obtain the following expression:
$\begin{split}\frac{\partial L_{LPO}\left(x_{1},x_{2}\right)}{\partial x_{1}}\bigg{/}\frac{\partial L_{LPO}\left(x_{1},x_{2}\right)}{\partial x_{2}}=\frac{-2\beta\cdot\text{sgn}(x_{1}-x_{2}-\frac{1}{2\beta})+C}{2\beta\cdot\text{sgn}(x_{1}-x_{2}-\frac{1}{2\beta})}\end{split}$ (10)
This simplification reveals that the ratio of $x_{1}$ and $x_{2}$ becomes a constant, and the relative magnitude of their gradients can be controlled by $\beta$ and $\lambda$. Meanwhile, to more effectively control the descent rate of $x_{2}$, we utilize the Straight-Through Estimator (STE) as introduced in Esser et al. (2021). This technique allows us to enhance Eq. 1 by enabling gradient propagation through discrete operations while maintaining the ability to optimize continuous variables effectively. As following:
$\begin{split}\begin{cases}L_{LPO-ste}^{x_{1}}=r_{1}\cdot 2\beta\left|x_{1}-x_{2}.\text{detach}()-\frac{1}{2\beta}\right|+\lambda\cdot\max(0,-x_{1})\\[8.61108pt] L_{LPO-ste}^{x_{2}}=r_{2}\cdot 2\beta\left|x_{1}.\text{detach}()-x_{2}-\frac{1}{2\beta}\right|+\lambda\cdot\max(0,-x_{1}.\text{detach}())\end{cases}\end{split}$ (11)
By applying the STE, we can isolate the gradients of the chosen and rejected log-probabilities, allowing for separate adjustment of their descent rates. This separation enhances the flexibility of our optimization process, facilitating finer control over the learning dynamics and improving overall model performance.
Ultimately, the expression for LPO-ste can be formulated as follows:
@@
In this expression, $L_{x_{1}}^{LPO-STE}$ and $L_{x_{2}}^{LPO-STE}$ represent the losses corresponding to the chosen and rejected responses, respectively, while $r_{1}$ and $r_{2}$ are coefficients that control the descent rates for these two components. By using the STE, we ensure that gradients are effectively managed during the optimization process, allowing for improved performance in preference alignment tasks.
In the practical application of LPO-ste, $r_{1}$ is typically fixed at 1.0, while $r_{2}$ is adjusted within the range $[0.05,3.0]$. Fig. 1 illustrates the descent rate of the rejected responses and the ascent rate of the chosen responses under varying $r_{2}$ values. From Fig. 1, it is evident that adjusting the size of $r_{2}$ facilitates control over the ascent rate of the chosen responses and the descent rate of the rejected responses. This capability enables us to fine-tune the model’s performance effectively.